OpenStudy (anonymous):

2x^2=10 does this problem have x intercepts

OpenStudy (anonymous):

You'll make y=0 to find x intercepts y=2x^2-10 2x^2-10=0 x^2-5=0 x^2=5 x1=sqrt5 x2=-sqrt5 x intercepts are (sqrt5,0);(-sqrt5,0)

OpenStudy (anonymous):

and the solution was 5,-5 right

OpenStudy (anonymous):

the correct solutions are-sqrt 5 and sqrt 5